# Convergence rates for polynomial optimization on set products

### v2

@@ 4.1 Problem statement and main result
Our final result, given in Theorem 12 below, provides the convergence rate of the moment hierarchy (39) for the order $2$ quantum Wasserstein distance, assuming that there is a bounded dual optimal solution of (38). This result is obtained by combining Lemma 10 with our analysis derived in Theorem 3. Let us denote by $\rho_{\min}$ and $\rho_{\max}$ the minimal and maximal respective eigenvalues of $\rho$, respectively. We define similarly $\nu_{\min}$ and $\nu_{\max}$. To bound the dual solution of (38), we will assume the following.
Assumption 11.
Both $\rho$ and $\nu$ have full rank and either $\nu_{\min}>\rho_{\max}$ or $\rho_{\min}>\nu_{\max}$.
@@
Let Assumption 11 hold. For any $t\geq 32n$, one has
$0\leq W^{2}_{2}(\rho,\nu)-W^{2}_{2}(\rho,\nu)_{2t}\leq\frac{\kappa(n)}{t^{2}},$ (42)
where $\kappa(n)$ is a constant depending only on $n$, $\rho$ and $\nu$, and being proportional to the constant $C_{\mathbf{X}}(2n,4)$ from Theorem 3. See relation (50) for details.
The proof of Theorem 12 is postponed in Section 4.3. Note that the hypotheses of Assumption 11 on $\rho$ and $\nu$ are rather restrictive. One can instead simply assume that there is an optimal solution $w^{\operatorname{opt}}=(\Lambda,\Gamma)$ for the dual (38) with bounded entries, in which case the constant $\kappa(n)$ is given as in (4.3).
@@ 4.3 Proof of Theorem 12
2. Under Assumption 11, let us handle specifically the case where $\nu_{\min}>\rho_{\max}$, the other case can be treated analogously. Let us show that in this case the second quantity $\|w^{\operatorname{opt}}\|_{1}$ is less than $2\sqrt{2}n\left(1-\frac{\rho_{\max}^{2}}{\nu_{\min}^{2}}\right)$. Note that every optimal tuple $(\Lambda_{1},\Lambda_{2},\Gamma_{1},\Gamma_{2})$ for the real dual (48) corresponds to an optimal tuple $(\Lambda,\Gamma)$ for the complex dual (38), with $\Lambda=\Lambda_{1}+\mathbf{i}\,\Lambda_{2}$ and $\Gamma=\Gamma_{1}+\mathbf{i}\,\Gamma_{2}$.
For all $z=a+\mathbf{i}\,b$ with $a,b\in\mathbb{R}$, one has $|a|+|b|\leq\sqrt{2}\sqrt{a^{2}+b^{2}}=\sqrt{2}|z|$, thus $\|\Lambda_{1}\|_{1}+\|\Lambda_{2}\|_{1}\leq\sqrt{2}\|\Lambda\|_{1}$, and $\|\Gamma_{1}\|_{1}+\|\Gamma_{2}\|_{1}\leq\sqrt{2}\|\Gamma\|_{1}$. For all $\delta>0$, the tuple $(\Lambda+\delta I,\Gamma-\delta I)$ is also optimal for (38) since
$\operatorname{Tr}(\rho(\Lambda+\delta I)+\nu(\Gamma-\delta I))=\operatorname{Tr}(\rho\Lambda+\nu\Gamma)+\delta\operatorname{Tr}(\rho-\nu))=\operatorname{Tr}(\rho\Lambda+\nu\Gamma),$
and
$\mathbf{x}^{*}(\Lambda+\delta)I\mathbf{x}-\mathbf{y}^{*}(\Gamma-\delta I)\mathbf{y}=\mathbf{x}^{*}\Lambda\mathbf{x}-\mathbf{y}^{*}\Gamma\mathbf{y},$
for all $(\mathbf{x},\mathbf{y})\in\mathbf{X}$. Thus, we assume without loss of generality that $\Lambda,-\Gamma\succeq 0$. Let us note $\mathbf{e}_{i}$ the vector having only 0 components except a component 1 in the $i$-th place. Since the polynomial $f-\mathbf{x}^{*}\Lambda\mathbf{x}-\mathbf{y}^{*}\Gamma\mathbf{y}$ is nonnegative on $\mathbf{X}$, one has in particular for $\mathbf{x}=\mathbf{y}=\mathbf{e}_{i}$:
$\Lambda_{ii}+\Gamma_{ii}\leq f_{\max}=2,$
implying that $\operatorname{Tr}(\Lambda)\leq 2-\operatorname{Tr}(\Gamma)$. In addition, the objective function $\operatorname{Tr}(\rho\Lambda)+\operatorname{Tr}(\nu\Gamma)$ is always nonnegative, thus
$\rho_{\max}\operatorname{Tr}(\Lambda)\geq\operatorname{Tr}(\rho\Lambda)\geq-\operatorname{Tr}(\nu\Gamma)\geq-\nu_{\min}\operatorname{Tr}(\Gamma).$
After combining the two above inequalities we obtain $\operatorname{Tr}(\Lambda)-\operatorname{Tr}(\Gamma)\leq 2\left(1-\frac{\rho_{\max}^{2}}{\nu_{\min}^{2}}\right)$.
The desired bound follows then from the Cauchy-Schwartz inequality:
$\|\Lambda_{1}\|_{1}+\|\Lambda_{2}\|_{1}+\|\Gamma_{1}\|_{1}+\|\Gamma_{2}\|_{1}$ $\leq\sqrt{2}\left(\|\Lambda\|_{1}+\|\Gamma\|_{1}\right)$
$\leq\sqrt{2}\sqrt{n}\left(\sqrt{\operatorname{Tr}(\Lambda^{2})}+\sqrt{\operatorname{Tr}(\Gamma^{2})}\right)$
$\leq\sqrt{2}n(\Lambda_{\max}-\Gamma_{\min})\leq\sqrt{2}n(\operatorname{Tr}(\Lambda)-\operatorname{Tr}(\Gamma)).$
where $\Lambda_{\max}$ and $\Gamma_{\min}$ denotes the largest and smallest eigenvalues of $\Lambda$ and $\Gamma$, respectively.
@@
$\kappa(n)\leq\frac{3}{2}\left(3+4\sqrt{2}n\left(1-\frac{\rho_{\max}^{2}}{\nu_{\min}^{2}}\right)\right)C_{\mathbf{X}_{\operatorname{Re}}(2n,4)}.$ (50)

### v3

@@ 4.1 Problem statement and main result
Our final result, given in Theorem 11 below, provides the convergence rate of the moment hierarchy (39) for the order $2$ quantum Wasserstein distance, assuming that there is a bounded dual optimal solution of (38). This result is obtained by combining Lemma 10 with our analysis derived in Theorem 3.
@@
Let assume that the dual (38) is attained at a bounded tuple. For any $t\geq 32n$, one has
$0\leq W^{2}_{2}(\rho,\nu)-W^{2}_{2}(\rho,\nu)_{2t}\leq\frac{\kappa(n,\rho,\nu)}{t^{2}},$ (42)
where $\kappa(n,\rho,\nu)$ is a constant depending only on $n$, $\rho$ and $\nu$, and being proportional to the constant $C_{\mathbf{X}}(2n,4)$ from Theorem 3. See relation (50) for details.
The proof of Theorem 11 is postponed in Section 4.3.
@@ 4.3 Proof of Theorem 11
2. Since the complex dual (38) is attained at a bounded tuple $(\Lambda,\Gamma)$, the real dual (48) is also attained at a bounded tuple $(\Lambda_{1},\Lambda_{2},\Gamma_{1},\Gamma_{2})$, with $\Lambda=\Lambda_{1}+\mathbf{i}\,\Lambda_{2}$ and $\Gamma=\Gamma_{1}+\mathbf{i}\,\Gamma_{2}$.
For all $z=a+\mathbf{i}\,b$ with $a,b\in\mathbb{R}$, one has $|a|+|b|\leq\sqrt{2}\sqrt{a^{2}+b^{2}}=\sqrt{2}|z|$, thus $\|\Lambda_{1}\|_{1}+\|\Lambda_{2}\|_{1}\leq\sqrt{2}\|\Lambda\|_{1}$, and $\|\Gamma_{1}\|_{1}+\|\Gamma_{2}\|_{1}\leq\sqrt{2}\|\Gamma\|_{1}$.
@@
$\kappa(n)\leq\frac{3}{2}\left(3+\sqrt{2}(\|\Lambda\|_{1}+\|\Gamma\|_{1})\right)C_{\mathbf{X}_{\operatorname{Re}}(2n,4)}.$ (50)